the water flows straight downward out at the end of circular pipe of radius r=2 mm and velocity =1m/s. it flows in a continuous column, without breaking into drops. What is the flow rate in this column?
1
Expert's answer
2021-09-12T19:00:06-0400

Given:

"r=2\:\rm mm=2*10^{-3}\: m"

"v=1\:\rm m/s"


The area of a pipe

"A=\pi r^2=3.14*(2*10^{-3}\:\rm m)=1.26*10^{-5}\:\rm m^2"

The flow rate

"Q=\frac{dV}{dt}=Av\\
=1.26*10^{-5}\:\rm m^2*1\: m/s=1.26*10^{-5}\:\rm m^3/s"